OpenStudy (mpj4):

isn't the integral of dx / [x (7x^2 - 5)^(1/2) ] = 1/a arcsec u/a??

OpenStudy (mpj4):

I ask because I applied it and I got it wrong.

ganeshie8 (ganeshie8):

looks you forgot to put the absolute value bars

OpenStudy (mpj4):

maybe

OpenStudy (mpj4):

my solution

OpenStudy (mpj4):

forgot to square the a=5 and the integral signs. Sorry, was a bit hasty making it.

OpenStudy (michele_laino):

you can try this substitution: \[\large \sqrt {7{x^2} - 5} = \sqrt 7 x + t\] where t is the new variable of integration

OpenStudy (mpj4):

Ok, thanks!